Suppose that 10 years ago you bought a home for $110,000, paying 10% as a down payment, and financing the rest at 9% interest for 30 years.
How much did you pay as a down payment?
How much was your existing loan for?
What is your current monthly payment on your existing mortgage?
How much total interest will you pay over the life of the existing loan?
1
Expert's answer
2018-02-15T08:21:07-0500
110000$*10%=11000$ -I paid as a down payment
110000$-11000$=99000$ was your existing loan for
Annuity payment can counting with formula
x=S×(P+P/((1+P)^N-1)), where x - monthly payment, S - existing loan, P-the monthly interest rate/100, N - number of months
x= 99000×(0.0075+0.0075/((1+0.0075^360-1))=796$
796×360=286560 - all payment
286560-99000=187560 - total interest will you pay over the life of the existing loan
The differentiated payment can counting with formula
on your existing mortgage is 99000/ 30 year × 20 year = 66000
66000/(20×12)=275$ - payment body of the loan
66000×9%/12=495 interest
275+495=570$ my payment passing ten year after buying per month